Compute Regression Analysis for following relationship: The relationship between death rate X1 vs. doctor availability X2. Doctor availability as a Predictor X, then death rate as a Response variable Y. Get the Regression Output and scatter plot between the variables using data analysis toolpak in Excel. X1= Death rate per 1000 residents. X2= doctor availability per 100,000 residents.

X1   X2
8 78
9.3 68
7.5 70
8.9 96
10.2 74
8.3 111
8.8 77
8.8 168
10.7 82
11.7 89
8.5 149
8.3 60
8.2 96
7.9 83
10.3 130
7.4 145
9.6 112
9.3 131
10.6   80
9.7 130
11.6 140
8.1 154
9.8 118
7.4 94
9.4 119
11.2 153
9.1 116
10.5 97
11.9 176
8.4 75
5 134
9.8 161
9.8 111
10.8 114
10.1 142
10.9 238
9.2 78
8.3 196
7.3 125
9.4 82
9.4 125
9.8 129
3.6 84
8.4 183
10.8 119
10.1 180
9 82
10 71
11.3 118
11.3 121
12.8   68
10 112
6.7 109
Answer #1

solution : X1 = Y........death rate, X2=X............doctor availability

\bar Y=\frac{\sum Y}{n}=\frac{493}{53}=9.31

\bar X=\frac{\sum X}{n}=\frac{6153}{53}=116.09

compute the following

dx=x-\bar X and   dy=y-\bar Y

regression equation of y on x =Y-\bar Y=\frac{\sum dxdy}{\sum dx^{2}}(X-\bar X)

Y-9.31=\frac{379.17}{ 74640.53}(X-116.09)

Y=0.005X-9.9

dxdy dy dx 1 8781.3138.09 49.8979 1450.85 2.59 59.91 155.1669 3589.21 2 9.3680.01 48.09 0.4809 2312.65 0.9141.09 37.3919 1688